OpenStudy (anonymous):

How to find the surface area of a cylinder with a diameter of 6ft and a height of 20cm using 3.14 rounded to the nearest whole number. (I cant figure it out my book doesn't give an example of it just a radius formula and ive tried searching the formula but i cant find any websites that show me how to do it or whats the formula)

OpenStudy (whpalmer4):

Well, the radius is just half of the diameter, right?

OpenStudy (whpalmer4):

Surface area of the cylinder is the area of the two ends + the area of the sides. The two ends are circles of radius \(r\):\[A_{ends} = 2*\pi r^2\] (the 2* is because there are two of them) The surface area of the side is found by slicing the cylinder down the side and unrolling it into a rectangle of height \(h\) and width \(2\pi r\) (which you should recognize as the circumference of the circle that makes up the ends). \[A_{sides} = 2\pi r h\](just length * width) \[A = 2\pi r^2 + 2\pi r h\]for the whole surface of the cylinder. Finally, \[r = \frac{d}{2}\] If you really want a formula for surface area based on diameter, you could substitute \(r = \frac{d}{2}\) into that equation: \[A = 2\pi (\frac{d}{2})^2 + 2\pi(\frac{d}{2})h = \pi\frac{d^2}{2} + \pi d h\]